A car was moving with avelocity of 30 m/s. the driver braked and the car stopped after 2 seconds.assuming that the acceleration of the car was constant during the pressure period , find the acceleration and the distance travelled by the car between the time the driver pressed the brakes and the stopping time

A car  was  moving with avelocity of 30 m/s. the driver braked and the car stopped after 2 seconds.assuming that the acceleration of the car was constant during the pressure period , find the acceleration and the distance travelled by the car between the time the driver pressed the brakes and the stopping time

Grade:12

2 Answers

vikas askiitian expert
13 years ago

initial velocity  = u = 30m/s

final velocity = v= 0

accleration = a     

we have , v = u + at

               0 = 30 + 2a               

  or      a = -30/2 = -15m/s2            (retardation)

now  , we have 

            v2  =  u2  +  2as

 v=0 , u =30 , a=-15

        0 = 302 + 2*(-15)s

       s = 30m                 

retardation = -15m/s2   &  distance covered before stopping is 30m

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Student

As acceleration is constant

Apply V=u+at

here final velocity V=0

a=-u/t = -30/2=-15 m/s2

-ve sign denotes deceleration (acceleration opposite to the direction of initial velocity)

now apply S=ut + at2/2

S=30*2 -15*22/2 = 60-30= 30m

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..